OpenStudy (anonymous):

x+(2z-4)-5z=1 solve for x

OpenStudy (anonymous):

just subtract everything one at a time except for the x x+(2z-4)-5z=1 x+(2z-4) = 1+5z x=1+5z-(2z-4) x=1+3z+4 x=3z+5

OpenStudy (anonymous):

thank you!

OpenStudy (skullpatrol):

Did you learn anything from that? @barbiegirl4513

OpenStudy (anonymous):

yes I just wasnt sure I could combine the ones that were in the ( )'s